OpenStudy (anonymous):

whoever gets this right gets a medal! How many combinations of 3 students can a teacher choose from 32 students?

OpenStudy (anonymous):

@karategirl2002 @@ganeshie8 @makako1018

OpenStudy (anonymous):

3/32?

OpenStudy (anonymous):

what? @jessienicole<3

OpenStudy (anonymous):

@jessienicole<3

OpenStudy (anonymous):

3 divided by 32?

OpenStudy (anonymous):

what does that equal? my calculator says 0.09375 @jessienicole<3

OpenStudy (anonymous):

3P32= 3P32/32= 3*2*1 32*31*30*29

OpenStudy (anonymous):

and so on

OpenStudy (anonymous):

what does that equal?! @moonlight101

OpenStudy (anonymous):

just give me the answer and I will be your fan and give you the medal @moonlight101

OpenStudy (anonymous):

it says im typing when im not so just ignore it

OpenStudy (anonymous):

\[32C3=\frac{32!}{3!(32-3)!}\]

OpenStudy (anonymous):

hello? I need help now!

OpenStudy (anonymous):

@moonlight101 @moonlight101

OpenStudy (anonymous):

I don't care about medals, but just to let you know: I am the only here to have given you the correct answer.

OpenStudy (anonymous):

29760

OpenStudy (anonymous):

i think

OpenStudy (anonymous):

Combinations, not permutations, @moonlight101

OpenStudy (anonymous):

o yeah dang it... idk the answer

OpenStudy (anonymous):

Scroll up. I already posted the answer.

OpenStudy (anonymous):

o haha :)